OpenStudy (anonymous):

A rectangle is inscribed with its base on the x-axis and its upper corners ont the parabola y=5-x^2

OpenStudy (anonymous):

\[y=5-x^2,x^2=-y+5=-\left( y-5 \right)\] this is a downward parabola with vertex at (0,5) cutting y-axis at (0,5) and x-axis at \[\left( -\sqrt{5},0 \right) ~and~\left( \sqrt{5},0 \right)\]  what you want to do,you can do now.